# Macroscopic entanglement between localized domain walls inside a cavity

### v1

@@ Appendix D Output Optical Spectrum
For this, here we provide a simple compact matrix method to calculate the spectrum for any quadrature set. The linearized quantum Langevin equations can be written in matrix form and expressed in the frequency domain:
@@
All the expectation values can be combined in the form of a matrix $\chi$: $\langle u^{\rm{in}}_{k}(\omega)u^{\rm{in}}_{l}(\omega^{\prime})\rangle=2\pi\chi_{k,l}(\omega)\delta(\omega+\omega^{\prime})$. The expression of spectral function then simplifies to
@@
The optical output spectrum is then obtained by summing over both the position and momentum cavity quadrature correlations $S(\omega)=S_{x_{a},x_{a}}(\omega)+S_{p_{a},p_{a}}(\omega)$.

### v2

@@ Appendix F Output optical spectrum
For this, here we provide a simple compact matrix method to calculate the spectrum for any quadrature set. The linearized quantum Langevin equations can be written as a matrix and expressed in the frequency domain:
@@
All the expectation values can be combined in the form of a matrix $\chi$: $\langle u^{\rm{in}}_{k}(\omega)u^{\rm{in}}_{l}(\omega^{\prime})\rangle=2\pi\chi_{k,l}(\omega)\delta(\omega+\omega^{\prime})$. The matrix elements $\chi_{k,l}$ can be obtained by using the Fourier transformation of the input noise correlations and converting them into quadrature variables
$\langle x^{\rm{in}}_{k}(\omega)x^{\rm{in}}_{l}(\omega^{\prime})\rangle$ $=2\pi\delta(\omega+\omega^{\prime})\frac{(1+2n^{\rm{th}}_{k})}{2}\delta_{k,l},$
$\langle p^{\rm{in}}_{k}(\omega)p^{\rm{in}}_{l}(\omega^{\prime})\rangle$ $=2\pi\delta(\omega+\omega^{\prime})\frac{(1+2n^{\rm{th}}_{k})}{2}\delta_{k,l},$
$\langle x^{\rm{in}}_{k}(\omega)p^{\rm{in}}_{l}(\omega^{\prime})\rangle$ $=2\pi\delta(\omega+\omega^{\prime})\frac{i}{2}\delta_{k,l},$
$\langle p^{\rm{in}}_{k}(\omega)x^{\rm{in}}_{l}(\omega^{\prime})\rangle$ $=-2\pi\delta(\omega+\omega^{\prime})\frac{i}{2}\delta_{k,l},$ (42)
where $k,l\in\{a,1,2\}$. Using this we get
$\chi=\begin{pmatrix}\frac{1+2n^{\rm{th}}_{a}}{2}&\frac{i}{2}&0&0&0&0\\ \frac{-i}{2}&\frac{1+2n^{\rm{th}}_{a}}{2}&0&0&0&0\\ 0&0&\frac{1+2n^{\rm{th}}_{1}}{2}&\frac{i}{2}&0&0\\ 0&0&\frac{-i}{2}&\frac{1+2n^{\rm{th}}_{1}}{2}&0&0\\ 0&0&0&0&\frac{1+2n^{\rm{th}}_{2}}{2}&\frac{i}{2}\\ 0&0&0&0&\frac{-i}{2}&\frac{1+2n^{\rm{th}}_{2}}{2}\end{pmatrix}$, (43)
and the expression of the spectral function simplifies to
@@
The optical output spectrum is then obtained by summing over both the position and momentum cavity quadrature correlations. Since the spectral function consists of noise correlations and expectation of quadrature variables in Eq. 42, it must satisfy the relation $S_{x_{a},x_{a}}(\omega)+S_{p_{a},p_{a}}(\omega)\geq 1$, which arises from the uncertainty relation of the noise spectrum $S_{x_{a},x_{a}}(\omega)S_{p_{a},p_{a}}(\omega)\geq 1/4$ [90, 91]. However, to achieve better contrast while plotting the spectrum, we define $S(\omega)=S_{x_{a},x_{a}}(\omega)+S_{p_{a},p_{a}}(\omega)-1$, such that $S(\omega)\geq 0$. This is the quantity plotted as the output optical spectrum in Fig. 4.
Figure 7: Pump power as a function of effective coupling $G_{\rm{eff}}$ and detuning $\tilde{\Delta}$. The light green dashed line is the phase transition line where the DW-DW entanglement is maximum, and the star marks the onset of entanglement as $\tilde{\Delta}$ decreases below resonance.
Appendix G Optical driving frequency and power
We analyze the optical frequency and power of the driving field necessary to achieve the desired DW-DW entanglement. Figure 7 shows the pump power, given by the relation $\hbar\omega_{a}|\xi|^{2}/2\kappa_{a}$, required to reach the different steady state phases in Fig. 2. The onset of entanglement between the two domain walls along the transition line (shown by a star in Fig. 7) is achieved with only a few mWs of power. For maximum attainable values of DW-DW entanglement a pump power of around 100 mW, operating in the THz regime, is needed. Such lasers have now been routinely realized in experiments [92, 93, 94].
We note that our formalism does not necessarily require a specific optical frequency, as only the detuning relative to the cavity frequency is relevant. However, a good choice of optical frequency is important in the design of the chiral cavity, as it needs to allow the necessary free spectral range of the cavity and keep in mind the associated limitations of cavity size and dimensions of the ferromagnetic strips inside the cavity. Moreover, with lasers in the infrared to visible regime, the operation is possible with slightly larger power requirements.